# The special unitary groups $SU(2n)$ as framed manifolds

### v3

@@ 3. Proofs of Theorem and Proposition
where $\Phi_{E}$ denotes the trivialization of the stable tangent space of $S(E)$ derived by the induced framing $\mathscr{L}_{S}$ on $G/S$.
Now $D(z_{2k-1},zz_{2k})\in Q_{k}$ equals $D(z_{2k-1},zz_{2k})=R_{1;2k-1}(z_{2k-1},0)R_{1;2k}(zz_{2k},0)$ where $R_{1;2k-1}(z_{2k-1},0)$ and $R_{1;2k}(zz_{2k},0)$, respectively, constitute the product components of $P\subset G$ which belong to $q_{0,2k-1}^{-1}((1,0)_{R})$ and $q_{0,2k}^{-1}((1,0)_{R})$. Putting $\mathscr{L}$ on $\Phi_{E}$, from this we see that the restrictions of $\mathscr{L}$ to $Q_{k}$ and $P$ overlap and therefore the restriction of $\mathscr{L}$ to $T_{k}\subset G$ must be stably trivial. This tells us that if we let $\rho$ denote the realification of the standard complex representation $G\hookrightarrow GL(m,\mathbb{C})$, then $\Phi_{E}$ must be stably equivalent to $\mathscr{L}^{(n-1)\rho}$. Substituting this into the equation above we have the desired result. ∎
@@
Let $m=2n+1$. Let $C$ be as above. Then clearly $S\cap C=\{I_{2n+1}\}$, so $S\times C$ becomes a subgroup of $G$. Letting $r:G\to G/C$ and $\tilde{r}:G/C\to G/(S\times C)$ be the quotient maps, the pair $(r,\tilde{r})$ gives a bundle map between principal $S$-bundles $G\to G/S$ and $G/C\to G/(S\times C)$. Consider the composition $r\circ\iota:\tilde{P}\hookrightarrow G\to G/C$ and $\tilde{r}\circ\psi:(S^{2})^{(m^{2}-m)/2}\times T^{n-1}\to G/S\to G/(S\times C)$ where $\iota$ denotes the inclusion of $\tilde{P}$ into $G$. Then clearly $(\tilde{r}\circ\psi)\circ\tilde{q}=\tilde{r}\circ(r\circ\iota)$ holds and $\dim((S^{2})^{(m^{2}-m)/2}\times T^{n-1})=\dim G/(S\times C)$. Following the proof of Lemma 3 we can see that $\tilde{r}\circ\psi$ is injective and so it becomes a homeomorphism. Therefore by replacing $\psi$ by $\tilde{r}\circ\psi$ in the proof of the theorem above we can get a modified form in the case of $m=2n+1$. This proves the proposition. ∎

### v4

@@ 3. Proofs of Theorem and Proposition
where $\Phi_{E}$ denotes the trivialization of the stable tangent space of $S(E)$ derived by the framing $\mathscr{L}_{S}$ on $G/S$ induced by $\mathscr{L}$.
Now according to the above definition the element $(z_{2k-1},\zeta_{2k})\in T_{k}$ $(1\leq k\leq n-1)$ represents $p(D(z_{2k-1},\zeta_{2k}))$ in $G/S$ via $\psi_{k}$ where $p:G\to G/S$ and also we know that $D(z_{2k-1},\zeta_{2k})$ can be written as
$D(z_{2k-1},\zeta_{2k})=R_{1;2k-1}(z_{2k-1},0)R_{1;2k}(\zeta_{2k},0)$
and that
$p(R_{1;2k-1}(z_{2k-1},0)R_{1;2k}(\zeta_{2k},0))=\phi(\ast)$
where $\phi$ is as above and $*=((1,0)_{R},\ldots,(1,0)_{R})\in(S^{2})^{(m^{2}-m)/2}$. From these relations we find that $\Phi_{E}$ corresponds to $\mathscr{L}$ which is trivialized over $T_{k}\subset G$ for $11\leq k\leq n-$ 1, namely that if we let $\rho$ denote the realification of the standard complex representation $G\to GL(m,\mathbb{C})$, then $\Phi_{E}$ can be taken to be $\mathscr{L}^{(n-1)\rho}$. This proves the theorem. ∎
@@
Let $m=2n+1$. Let $C$ be as above. Then clearly $S\cap C=\{I_{2n+1}\}$, so $S\times C$ becomes a subgroup of $G$. Letting $r:G\to G/C$ and $\tilde{r}:G/C\to G/(S\times C)$ be the quotient maps, the pair $(r,\tilde{r})$ gives a bundle map between principal $S$-bundles $G\to G/S$ and $G/C\to G/(S\times C)$. Consider the composition $r\circ\iota:\tilde{P}\hookrightarrow G\to G/C$ and $\tilde{r}\circ\psi:(S^{2})^{(m^{2}-m)/2}\times T^{n-1}\to G/S\to G/(S\times C)$ where $\iota$ denotes the inclusion of $\tilde{P}$ into $G$. Then clearly $(\tilde{r}\circ\psi)\circ\tilde{q}=\tilde{r}\circ(r\circ\iota)$ holds and $\dim((S^{2})^{(m^{2}-m)/2}\times T^{n-1})=\dim G/(S\times C)$. Following the proof of Lemma 3 we can see that $\tilde{r}\circ\psi$ is a degree one map. Therefore by replacing $\psi$ by $\tilde{r}\circ\psi$ in the proof of the theorem above we can get a modified form in the case of $m=2n+1$. This proves the proposition. ∎
Remark .
From the proof of the theorem we see that
$e_{\mathbb{C}}([SU(2n),\mathscr{L}])=0.$
 [3]. The doubling of the framing occurred there can be dissolved by thinking of the restriction of $E$ to $T_{k}$ for every $k$ as a trivial complex line bundle. But instead its first Chern class becomes zero and so, according to Proposition 2.1 of  [6], the value of $e_{\mathbb{C}}$ must become zero.